OpenStudy (anonymous):

solve for x 9x + 2= 4x + 32

OpenStudy (anonymous):

9x + 2= 4x + 32 9x-4x=32-2 5x=30 x=6

OpenStudy (anonymous):

5x=30 x=6

OpenStudy (anonymous):

Change the place of 4x to have X s together : 9x-4x=32-2 and numbers too ! @cassytaylor92 and @tmac are very right ! thanks ;) And welcome @tmac

OpenStudy (anonymous):

thanks! I just found this site...what fun!

OpenStudy (anonymous):

;)